Find the maximum value of the function f(x) = x + x2
1
Expert's answer
2019-09-24T11:19:25-0400

The derivative of "f(x) = x + x^2" is "f'(x) = 1 + 2 x". Equation "f'(x) = 0" has solution "x = -\frac{1}{2}", and the signs of the derivative change from negative to positive in the intervals "(-\infty, -\frac{1}{2})" , "(-\frac{1}{2}, \infty)" , hence "x = -1\/2" is a local minimum. The function is not bounded from above, hence its maximum value is "+\infty".